OpenStudy (anonymous):

Find the product of 2x^4(2x^2 + 3x + 4). 2x^8 + 3x^4 + 4x^4 4x^6 + 6x^5 + 8x^4 4x^4 + 3x^5 + 2x^6 3x^6 + 4x^5 + 5x^4

OpenStudy (shamil98):

You can easily find the answer without having to distribute it to all terms what is 2x^4 * 2x^2 ? 2 x 2 = ? x^4 * x^2 = Remember \[\large a^m \times a ^n = a^{m+n}\]